Agenda/Minutes of the VIS Meeting on 2016/11/4
2016/11/4 13:30 - 14:40
Participants: Takahashi, Okutomi, Tanaka, Mark, Aso, Ishizaki, Shoda, Sato, Fujii, Akutsu, Enzo
- Installation: The review was held on 4th.
- Pre-isolator: The new shaft screws of motor were delivered. They will be replaced.
- The transportation of the PI to the Y-end may be delayed by 1 week due to the site preparation schedule.
- EQ stop, release sensor, dummy payload: Production is on going.
- BF damper: Open bidding is on 9th.
- Maraging wire: New design was checked by CRY. Asking the company a quotation.
Type B (Mark)
- Type B Past Week Summary
Updated the BS circuit diagram D1503600:
- Added information on the binary IO.
- Renamed channels to conform to Okutomi-kun's new scheme.
- Made the ADC/DAC channel assignment consistent and more logical.
- Updated the stepper motor sections with info about the new driver chassis.
- Worked on schedule.
- Updated the installation procedure with lessons learned from the last few weeks.
- Downloaded Hirata-san's Inventor model of the BS in the tank and starting thinking about installation of the suspension and lower bread board.
- Mark and Hirata-san worked on collecting parts and tools that will be needed for cabling of the BS next week. (11 special-purpose cables need to be made for the picos/steppers/OSEMs/LVDT on the RM/IRM/IM/BF and, time permitting, 2 more for the SF.)
- Updated the BS 3D Assembly with the new rods and receptacles.
- Made 2D drawings of SR2 and SR3 (ongoing).
- Made SRM 3D assembly (ongoing)
- Asked for and got quotes for hexagon cable clamps. Takahashi-san ordered the PEEK components. Next week Hirata-san will look for parts that should already be at Kamioka.
- Washed screws for hexagon cable clamps and shim rings for Fabian.
- The ordering process finished. We told the company to proceed the manufacturing. The detailed discussion with the company will be help at 14th 11:00-.
- We have got the cost estimate. The ordering process is under discussion.
Flag removing jig
- The design is fixed. The material will be bought by us and will be passed ATC to machine.
The number of the cables used for Type-Bp is checked. Passed information to Takahashi-san.
The number is enough only for Type-Bp. I am not sure if Type-B and Type-A are included.
Also we have got the cost estimate for the magnets. They will be ordered soon.
The chain blocks and lever blocks will also be ordered soon.
- We have got the costestimate. We are asking for the competitive quotes.
- We are going to order the OB related parts necessary in only this FY for now.
- We are going to disassemble the iKAGRA PR3 from 9th. The mirror box and the perican box is sent to Kamioka.
- Put a report here
Watchdog system (Tanaka)
Design and preparation(investigation of procurement) for production.PDF
- OSEM signal is picked up through 3kOhm voltage divider on the hot line. This should be increased to 30kOhm. Also the same resistor should be connected to the cold line to keep the balance.
- The sensing resistor for the coil current is inserted in the hot line. Since this is the current driving circuit, the imbalance between the hot and cold will not be a problem.
Will show the design at the next AEL meeting.--->Nov.14th.2016 3pm(Schedule of presentation for WDS.)
- No issue
- First contact application to PR3
- Currently the ventilation system in the tunnel is not yet usable (at least for a month or so)
- We cannot apply first contact in the tunnel yet.
- We will send the PR3 back to Kashiwa without first contact
- How to apply the first contact on the HR surface of a mirror with flags attached.
- With the current design of the mirror box, we cannot access the HR surface. Some modification is necessary.
- Will have a discussion in the next week in Kamioka.
- Naming convention of the suspension stages
Marionetta -> MT
- How to call the last stage should be discussed at the chief meeting.
Site work plan for the next week
- Mon: Accelerometer test, Extended Chief Meeting
- Tue: BS Test Hanging, Accelerometer test, PR3 disassemble, Fire Drill
- Wed: BS Test Hanging, Accelerometer test, PR3 disassemble
- Thu: BS Test Hanging, PR3 disassemble
- Fri: BS Test Hanging, PR3 disassemble
- Travel (Week of 11/7):
- Takahashi: 7-9 (Extended chief meeting/Fire drill/Accelerometer test)
- Shoda: 7-11 (Extended chief meeting/Fire drill/PR3 disassemble)
- Mark: 8 - 11 (Fire drill/BS)
- Okutomi: 7 - 11 (Extended chief meeting/Fire drill/Accelerometer test/Realtime code)
- Hirata: 8 - 11 (Fire drill/BS)
- Sato: 7-9 (Fire drill/Accelerometer test)
- Enzo: 8-11 (PR3 disassembly)
- Travel (Week of 11/14):
VIS F2F: Nov. 18, Dec. 2